The density or of carbon dioxide is 1.98kg/m^3. express the density of carbon dioxide in grams/cm^3
1
Expert's answer
2020-01-22T04:18:06-0500

As 1kg = 1,000 g and 1m3=1,000,000cm3:

1.98 kg/m3 = 1.98 × 1,000 g / 1,000,000 cm3 = 1.98 × 10-3 g/cm3 = 0.00198 g/cm3.

Answer: 0.00198 g/cm3.